OpenStudy (anonymous):

magine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

I'd tell the student to do the same thing to both expressions to ensure they stay as an equation:\[2(x+6)=22+4(x-1)\]

OpenStudy (anonymous):

Then I'd say use the distributivity rule of brackets to expand:\[2x+12=22+4x-4\]

OpenStudy (anonymous):

1) Distribute the parenthesis using the Distributive property a(b-c)=ab-ac 2)Then using the addtion and subtraction property of equality, move all your x variables to one sine 3) using the division property of equality to get x by itself (x=?)

OpenStudy (anonymous):

Then I'd tell them to simplify, as in the first step:\[2x=6+4x\]

OpenStudy (anonymous):

Finally, get all the x on one side and constants on the other:\[2x=-6\]

OpenStudy (anonymous):

Then divide all by 2!

OpenStudy (anonymous):

2(x + 6) − 10 = 12 + 4(x − 1) 2x + 12 - 10 = 12 + 4x -4 look i have multiplied 2 and 4 with the bracket terms.. now.. 2x + 2 = 4x +12 - 4 2x +2 = 4x +8 subtract 2 on both sides 2x + 2 -2 = 4x + 8 -2 2x = 4x +6 now subtract 4x on both sides 2x-4x = 4x -4x +6 -2x = 6 divide both sides by -2 -2x/-2 = 6/-2 x= -3

OpenStudy (anonymous):

the addition and subtraction property of equality states that if you subtract or add something to one side of the equal sign, you must do the same to the other sides For example, \[1=1\] What you shouldn't do \[1+1=1\] \[2\neq 1\] What is Correct \[1+1=1+1\] \[2=2\]

OpenStudy (anonymous):

thank you guys for all the help